A crystal has a=2,b=3 and c=3oA and α=β=γ=90o. What is the length of the longest diagonal?
1
Expert's answer
2020-03-23T10:12:55-0400

Since a b and c are perpendicular then the length of longest diagonal will be equal to

"\sqrt{a^2+b^2+c^2}=\sqrt{2^2+3^2+3^2}=\sqrt{22} \ A^o"